OpenStudy (anonymous):

4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

please someone help!!!

OpenStudy (alekos):

just boils down to two equations x - 0.525 = .01 and x-0.525 =-.01

OpenStudy (anonymous):

k, I get that

OpenStudy (alekos):

|x| = a means x=a and x=-a

OpenStudy (anonymous):

I get that but I don't get how to solve it.

OpenStudy (welshfella):

just add 0.525 to both sides of the equations

OpenStudy (welshfella):

x - 0.525 = .01 +0.525 + 0.525 x = -01 + 0.525

OpenStudy (anonymous):

x-0.525=0.01 x-0.525= -0.01 +0.525 +0.525 x= 0.535 x= 0.515 Right?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

omergersh thank you.... I don't get this so yeah

OpenStudy (anonymous):

could you help me with one more?